# Time-dependent dynamics in the confined lattice Lorentz gas

### v1

@@ III.2 Obstructed biased motion
The only non-vanishing matrix elements for an obstacle located in the origin are the obstacle site $\textbf{e}_{0}=\textbf{0}$ and its nearest neighbors $\textbf{e}_{\pm 2}=\pm\textbf{e}_{y},\textbf{e}_{\pm 1}=\pm\textbf{e}_{x}$. Since the potential $\hat{v}$ displays non-vanishing contributions only in the distinguished basis $\textbf{r}\in\textbf{0}\cup\mathcal{N}$ it is sufficient to consider Eq. (29) in such a subspace only. The bare propagator reads
$G_{0}=\begin{pmatrix}g_{00}&g_{11}e^{F/2}&g_{01}&g_{11}e^{-F/2}&g_{02}\\ g_{11}e^{-F/2}&g_{00}&g_{10}e^{-F/2}&g_{20}e^{-F}&g_{11}e^{-F/2}\\ g_{01}&g_{10}e^{F/2}&g_{00}&g_{10}e^{-F/2}&g_{01}\\ g_{11}e^{F/2}&g_{20}e^{F}&g_{10}e^{F/2}&g_{00}&g_{11}e^{F/2}\\ g_{02}&g_{11}e^{F/2}&g_{01}&g_{11}e^{-F/2}&g_{00}\end{pmatrix}$ (30)
some comments are in order. The quantity $g_{lm}$ denotes the propagator for $l$ steps along the $x$ direction and $m$ steps along the $y$ direction. The free propagators are defined via $g_{10}(s):=\langle\textbf{e}_{x}|\hat{G}_{0}(s^{\prime})|\textbf{0}\rangle$, $g_{11}(s):=\langle\textbf{e}_{x}|\hat{G}_{0}(s^{\prime})|\textbf{e}_{y}\rangle$, $g_{01}(s):=\langle\textbf{0}|\hat{G}_{0}(s^{\prime})|\textbf{e}_{y}\rangle$, $g_{02}(s):=\langle\textbf{0}|\hat{G}_{0}(s^{\prime})|2\textbf{e}_{y}\rangle=\langle\textbf{e}_{y}|\hat{G}_{0}(s^{\prime})|\!-\!\textbf{e}_{y}\rangle$. Note that the entries of the matrix $G_{0}$ are the propagators in the presence of the external force $F$. As we showed in Appendix A, the presence of the bias amounts to the factor $e^{\pm F/2}$ in front of the free propagator which is computed at the shifted frequency $s^{\prime}\equiv s+\Gamma-1$. Therefore, e.g., $g_{00}$ in [Eq. (30)] stands for $g_{00}(s^{\prime})$. The quasi-confinement along $y$ implies that the transition amplitudes for $n$ steps along $x$ differs from the corresponding one for $n$ steps along $y$, e.g., $g_{10}\neq g_{01}$ and the equality is restored only in the limit $L\rightarrow\infty$.
The solution of the scattering problem [Eq. (29)] amounts to solve a $5\times 5$ matrix problem, something that can be conveniently pursued with the aid of symbolic computation programs. At a formal level we can write the solution by introducing the $S$-matrix,
$S=(1-vG_{0})^{-1}\,,$ (31)
which allows us to express the forward scattering matrix as follows
$t=Sv\,.$ (32)
@@ IV Velocity relaxation
This result has been obtained by employing the symmetry-adapted basis outlined in Appendix C which is inspired by the problem in absence of driving [36]. Although some of the symmetries are lost due to the driving, the residual symmetries are conveniently handled by the orthogonal transformation used in absence of driving. The right hand side of Eq. (34) is expressed in terms of a matrix element involving the forward scattering $t$-matrix. The vector $|n\rangle$ is related to conservation of probability via $\langle n|\hat{v}=0$, it is therefore a "neutral mode”. The vector $|p_{x}\rangle$ is reminiscent of a “dipole” oriented along the $x$-axis. In general, the matrix $t(\textbf{k})$ in plane wave basis can be decomposed into different channels corresponding to the aforementioned modes, as well as the dipole $|p_{y}\rangle$, the quadrupole $|d_{xy}\rangle$, and the $s$-wave $|s\rangle$. Rather elegantly, for the calculation of the velocity response the only non-vanishing contribution is the one that couples the neutral mode to the $p_{x}$-channel.
@@
The analytic expression of the function $\mathcal{V}_{L}(F;s)$ turns out to be rather cumbersome to be reported here, however, in Appendix C.2, we illustrate how to obtain this function. A plot of $\mathcal{V}_{L}(F;0)$ as a function of $F$ is provided in Fig. 3 of the accompanying Letter [26].
@@
It is meaningful to ask what is the role played by the strength of the driving for a certain confinement $L$. This is addressed in the plot of Fig. 3 in which the normalized velocity response is plotted in double-logarithmic scale for several values of the force at $L=2$ and, for comparison, also $L=\infty$. Quite interestingly, as $F\rightarrow 0$ the velocity approaches the asymptotic value with a power-law of the form $t^{-d_{E}/2}$ with the effective dimension $d_{E}=1$, this is shown in Fig. 3 with a dashed black line for $L=2$, however this feature persists for any finite $L$. For comparison we also illustrate the relaxation for the unbounded model, in this case the power-law is characterized by $d_{E}=2$, as shown with the solid black line in Fig. 3.
These different exponents can be obtained as follows: within linear response – legitimate for the regime $F\rightarrow 0$ – the tail $t^{-d_{E}/2}$ is inherited from the long-time decay of the VACF; in particular, the velocity approach is given by $\delta v(t)=-F\int_{t}^{\infty}Z(t^{\prime})\textrm{d}t^{\prime}$. From Fig. 3 one can infer that even for a very small driving the relaxation is no longer a power-law. In fact, the power-law tail is termed “fragile” in the sense that an infinitesimally small driving entails a decoration of the power-law with an exponential decay in the external force, as shown in Fig. 3 and in the inset of Fig. 2. Hence, the drift $v(t)$ approaches the stationary state $v(t\rightarrow\infty)$ exponentially fast even for an arbitrarily small but finite driving force $F$, a result in strong contradistinction to the FDT predicting a slow algebraic approach.
@@ V.1 Velocity autocorrelation function and diffusion coefficient at equilibrium
For any finite $L$ the prefactor $A_{L}$ is finite while for $L\rightarrow\infty$ we have $A_{L}\sim p/L$ with coefficient $p=\pi^{3/2}/8$. In the accompanying Letter we show that for large but finite $L$ there exists a transient time domain in which the decay is of the form $\propto t^{-2}$ and that the long-time tail $\propto t^{-3/2}$ is ultimately attained for $t\gg t_{L}$ where $t_{L}\propto L^{2}/D$ is the time scale associated to the exploration of the periodic direction. Hence, we plot the VACF rescaled to its asymptotic tail [Eq. (55)] as a function of the rescaled time $\tilde{t}:=t/t_{L}$; without loss of generality we can take $t_{L}\equiv L^{2}$ and the result we obtain is the shown in Fig. 4. We observe that the data collapse to the master curve protracts from large scaled times $\tilde{t}\gg 1$ down to small times $\tilde{t}\ll 1$ upon increasing $L$. We further note that the scaling breaks down for microscopic times $t=O(\tau)$; in terms of rescaled times this implies a behavior proportional to $\tilde{t}^{-2}$ for $\tilde{t}\gg\tau/t_{L}$ and $\tilde{t}^{-3/2}$ for $\tilde{t}\gg 1$. In the limit $L\rightarrow\infty$ and $t\rightarrow\infty$ with fixed $t/t_{L}$ the data exhibit a collapse onto a single master curve whose behavior at large times is $\tilde{t}^{-3/2}$ while for short times it behaves as $\pi^{-1/2}\tilde{t}^{-2}$, these asymptotic power laws are indicated with dashed and dot-dashed black curves in Fig. 4.
Figure 4: The VACF rescaled to its asymptotic tail $-nA_{L}t_{L}^{-3/2}$. The dashed black line indicates the exact long-time tail $\propto\tilde{t}^{-3/2}$, the dot-dashed corresponds to the transient tail $\propto\tilde{t}^{-2}$.
The black curve in Fig. 4 is obtained by inverting the Laplace transform $\widehat{Z}_{L}(s)$ for the unbounded model. The latter is formally obtained by taking the limit $L\rightarrow\infty$ in $\widehat{Z}_{L}(s)=[1+n(1+\mathcal{V}_{L}(0,s))]/4$, where $\mathcal{V}_{L}(0;s)=-8/\Delta_{L}(s)$ is the function already discussed in the context of the velocity relaxation, however, in this case it is evaluated for $F=0$ and $s\neq 0$. As anticipated, the limit can be performed at a formal level while its implementation is far from being trivial since propagators themselves are expressed as a sum over modes $k_{y}=2\pi q/L$ labelled by an integer running from $q=1$ to $q=L-1$; we refer to Appendix A for further details on lattice Green’s functions. Nonetheless, results for $L\rightarrow\infty$ are more conveniently retrieved simply by examining the solution of the unbounded model which yields [27, 34], in our notations,
@@
where $\textsf{E}[\kappa]=\int_{0}^{\pi/2}\textrm{d}\theta\left(1-\kappa\sin^{2}\theta\right)^{1/2}$ is the complete elliptic integral of the second kind [44, 45]; in this paper we adopted Mathematica’s convention for elliptic integrals. For the sake of completeness we also provide the quantity $\Delta_{L}(s)$ for the unbounded model. The the matrix element of the forward scattering matrix can expressed in terms of an elliptic integral and the result takes the form
@@
the last line holds as $s\rightarrow 0$ and is useful for the analysis of the long-time tail. The small-frequency expansion of the VACF contains smooth terms in $s$ and the first non-regular term in such an expansion is proportional to $\propto s\ln s$ with prefactor $-\pi/8$. The latter yields the leading form of the long-time tail which, including the prefactor, is
@@ V.2 Diffusion coefficient and mean-square displacement away from equilibrium
where $\Xi_{L}(F)$ is known exactly for any system size $L$ and arbitrary force $F$. In particular, for $F=0$ one finds $\Xi_{L}(0)=(1/4)-2/C_{L}$ and the corresponding result in equilibrium [Eq. (48)] is retrieved. In order to single out the effects due to the disorder we consider the force-induced diffusion coefficient corrected by the empty lattice ($n=0$), this subtraction entails
@@
2 0.4723 $44\sqrt{2}-$ 4 $1.6569$
3 0.9556 $44\sqrt{7/3}-$ 4 $2.1101$
4 1.2412 $42\sqrt{2}+2\sqrt{3}-$ 4 $2.2925$
@@ VI Summary and conclusions
We find that – already in equilibrium – the quasi-confined lattice Lorentz model in the absence of the external force displays the intriguing feature of a dimensional crossover. We showed this property by an exact analysis of the equilibrium VACF. The latter is characterized by a transient long-time tail of the form $t^{-2}$ – reminiscent of the unbounded model –, while the long-time asymptotic behavior is a power law $\propto t^{-3/2}$. These different behaviors can be rationalized in terms of a dimensional-dependent long-time tail of the form $t^{-(d_{E}+2)/2}$, where $d_{E}$ is an effective dimensionality. The analytic solution shows that for any finite $L$ the effective dimensionality is $d_{E}=1$ at large times while $d_{E}=2$ at intermediate times. The two regimes are separated by a size-dependent time scale $t_{L}=L^{2}/D$ which can be interpreted as the time to explore the confined direction.
@@
We then switched on the driving force. When the tracer particle is pulled a non-equilibrium stationary state is reached in which the average motion occurs with a force-dependent terminal velocity that we found analytically. The instantaneous velocity relaxes towards its terminal value via a power-law $t^{-d_{E}/2}$ decorated with an exponential time-dependence due to the force; the exponent is $d_{E}=1$ for any finite confinement $L$. As a result, the long-time tail is fragile, as it happens for the unbounded model and in such a case $d_{E}=2$. We analyzed the fluctuations of the tracer in the direction parallel to the force. For large times the variance of the displacement along the force grows linearly with time and its prefactor is a force-dependent and confinement-dependent diffusion coefficient. For large forces the confinement plays a minor role, as expected on intuitive basis since the particle hops rarely in the confined direction. On the other hand the diffusion coefficient turns out to be sensitive to the confinement in the regime of small driving. In this regime the force-dependent diffusion coefficient of the unbounded model exhibits non-analytic behavior of the form $F^{2}\ln|F|$. The occurrence of non-analyticity persists upon confinement, now with a non-analytical behavior of the form $b_{L}|F|$, with an amplitude $b_{L}$ that we obtained analytically. Although it is expected the suppression of diffusivity upon increasing the disorder, we found that this is true only when the force is smaller than a critical force $F_{c,L}$. Quite interestingly, if $F>F_{c,L}$ then the increase of disorder leads to an enhanced diffusivity. We find that $F_{c,L}$ decreases upon increasing $L$, meaning that the onset for the disorder-enhancement diffusion regime is favored by the confinement. Finally, the complete time-dependent solution shows that the behaviors at short and large times are interpolated by a transient superdiffusive behavior. We characterized this feature in terms of the local exponent of anomalous diffusion $\alpha(t)$ by showing how it increases upon increasing of the force. In this regard, especially for large forces, the confined model behaves analogously to the unbounded one and therefore the saturation to $\alpha(t)=3$ is expected as a transient phenomenon followed by diffusion at infinite times.

### v2

@@ III.2 Obstructed biased motion
The only non-vanishing matrix elements for an obstacle located in the origin are the obstacle site $\textbf{e}_{0}=\textbf{0}$ and its nearest neighbors $\textbf{e}_{\pm 2}=\pm\textbf{e}_{y},\textbf{e}_{\pm 1}=\pm\textbf{e}_{x}$. To be definite, the distinguished basis is $\mathcal{B}=(-\textbf{e}_{y},-\textbf{e}_{x},\textbf{e}_{0},\textbf{e}_{x},\textbf{e}_{y})$; see Fig. 2.
Figure 2: The distinguished basis $\mathcal{B}$ centered on the impurity.
Since the potential $\hat{v}$ displays non-vanishing contributions only in the distinguished basis $\textbf{r}\in\textbf{0}\cup\mathcal{N}$ it is sufficient to consider Eq. (29) in such a subspace only. The bare propagator reads
$G_{0}=\hskip-4.2679pt\begin{pmatrix}g_{00}&g_{11}e^{F/2}&g_{01}&g_{11}e^{-F/2}&g_{02}\\ g_{11}e^{-F/2}&g_{00}&g_{10}e^{-F/2}&g_{20}e^{-F}&g_{11}e^{-F/2}\\ g_{01}&g_{10}e^{F/2}&g_{00}&g_{10}e^{-F/2}&g_{01}\\ g_{11}e^{F/2}&g_{20}e^{F}&g_{10}e^{F/2}&g_{00}&g_{11}e^{F/2}\\ g_{02}&g_{11}e^{F/2}&g_{01}&g_{11}e^{-F/2}&g_{00}\end{pmatrix}$. (30)
Some comments are in order. The quantity $g_{lm}$ denotes the propagator for $l$ steps along the $x$ direction and $m$ steps along the $y$ direction. The entries of this matrix can be readily obtained by referring to the distinguished basis illustrated in Fig. 2. For the sake of clarity, we illustrate how to find them through some examples. The element $(1,3)$ represents the jump from $\textbf{e}_{0}$ to $\textbf{e}_{y}$, therefore it corresponds to the propagator with $l=0$ and $m=1$, this yields $g_{01}$. Then, the entry $(4,1)$ corresponds to the jump from $\textbf{e}_{y}$ to $\textbf{e}_{x}$, therefore it is associated with $l=1$ and $m=1$ and since the jump is parallel to the force, the propagator is multiplied by $\exp(F/2)$, for the reasons proved in Appendix A, this gives $g_{11}\exp(F/2)$. The free propagators are defined via $g_{10}(s):=\langle\textbf{e}_{x}|\hat{G}_{0}(s^{\prime})|\textbf{0}\rangle$, $g_{11}(s):=\langle\textbf{e}_{x}|\hat{G}_{0}(s^{\prime})|\textbf{e}_{y}\rangle$, $g_{01}(s):=\langle\textbf{0}|\hat{G}_{0}(s^{\prime})|\textbf{e}_{y}\rangle$, $g_{02}(s):=\langle\textbf{0}|\hat{G}_{0}(s^{\prime})|2\textbf{e}_{y}\rangle=\langle\textbf{e}_{y}|\hat{G}_{0}(s^{\prime})|\!-\!\textbf{e}_{y}\rangle$. Note that the entries of the matrix $G_{0}$ are the propagators in the presence of the external force $F$. As we showed in Appendix A, the presence of the bias amounts to the factor $e^{\pm F/2}$ in front of the free propagator which is computed at the shifted frequency $s^{\prime}\equiv s+\Gamma-1$. Therefore, e.g., $g_{00}$ in [Eq. (30)] stands for $g_{00}(s^{\prime})$. The quasi-confinement along $y$ implies that the transition amplitudes for $n$ steps along $x$ differ from the corresponding one for $n$ steps along $y$, e.g., $g_{10}\neq g_{01}$ and the equality is restored only in the limit $L\rightarrow\infty$.
The solution of the scattering problem [Eq. (29)] amounts to solve a $5\times 5$ matrix problem, something that can be conveniently pursued with the aid of symbolic computation programs. At a formal level, the solution is
t=Sv,t=Sv\,, (31)
where the $S$-matrix is
$S=(1-vG_{0})^{-1}\,.$ (32)
@@ IV Velocity relaxation
This result has been obtained by employing the symmetry-adapted basis outlined in Appendix C which is inspired by the problem in absence of driving [56]. Although some of the symmetries are lost due to the driving, the residual symmetries are conveniently handled by the orthogonal transformation used in absence of driving. The right-hand side of Eq. (34) is expressed in terms of a matrix element involving the forward scattering $t$-matrix. The vector $|n\rangle$ is related to conservation of probability via $\langle n|\hat{v}=0$, it is therefore a "neutral mode”. The vector $|p_{x}\rangle$ is reminiscent of a “dipole” oriented along the $x$-axis. In general, the matrix $t(\textbf{k})$ in plane wave basis can be decomposed into different channels corresponding to the aforementioned modes, as well as the dipole $|p_{y}\rangle$, the quadrupole $|d_{xy}\rangle$, and the $s$-wave $|s\rangle$. Rather elegantly, for the calculation of the velocity response the only non-vanishing contribution is the one that couples the neutral mode to the $p_{x}$-channel.
@@
The analytic expression of the function $\mathcal{V}_{L}(F;s)$ turns out to be rather cumbersome to be reported here; nevertheless, in Appendix D we show the explicit result for the case $L=2$. The procedure to obtain $\mathcal{V}_{L}(F;s)$ is illustrated in Appendix C.2 and explicit results for this function are available for any $L$ in the supporting Mathematica script [62].
A plot of $\mathcal{V}_{L}(F;0)$ as a function of $F$ is provided in Fig. 3.
Figure 3: The function $\mathcal{V}_{L}(F;0)$ as a function of $F$ for some values of the quasi-confinement size $L$. Symbols show the results obtained from stochastic simulations.
@@
It is meaningful to ask what is the role played by the strength of the driving for a certain confinement $L$. This is addressed in the plot of Fig. 5 in which the normalized velocity response is plotted in double-logarithmic scale for several values of the force at $L=2$ and, for comparison, also $L=\infty$. Quite interestingly, as $F\rightarrow 0$ the velocity approaches the asymptotic value with a power-law of the form $t^{-1/2}$, this is shown in Fig. 5 with a dashed black line for $L=2$, however this feature persists for any finite $L$. For comparison we also illustrate the relaxation for the unbounded model, in this case the power-law is characterized by the power law $t^{-1}$, as shown with the solid black line in Fig. 5.
These different exponents can be obtained as follows: within linear response – legitimate for the regime $F\rightarrow 0$ – the tail $t^{-(d+2)/2}$ is inherited from the long-time decay of the VACF; in particular, the velocity approach is given by $\delta v(t)=-F\int_{t}^{\infty}Z(t^{\prime})\textrm{d}t^{\prime}$. From Fig. 5 one can infer that even for a very small driving the relaxation is no longer a power-law. In fact, the power-law tail is termed “fragile” in the sense that an infinitesimally small driving entails a decoration of the power-law with an exponential decay in the external force, as shown in Fig. 5 and in the inset of Fig. 4. Hence, the drift $v(t)$ approaches the stationary state $v(t\rightarrow\infty)$ exponentially fast even for an arbitrarily small but finite driving force $F$, a result in strong contradistinction to the FDT predicting a slow algebraic approach.
@@ V.1 Velocity autocorrelation function and diffusion coefficient at equilibrium
For any finite $L$ the prefactor $A_{L}$ is finite while for $L\rightarrow\infty$ we have $A_{L}\sim p/L$ with coefficient $p=\pi^{3/2}/8$. For the sake of completeness, the occurrence of the long-time tail $\propto t^{-3/2}$ has been reported for a one-dimensional Lorentz-type model [64] although in such a model the VACF exhibits an oscillatory behavior. In the accompanying Letter we show that for large but finite $L$ there exists a transient time domain in which the decay is of the form $\propto t^{-2}$ and that the long-time tail $\propto t^{-3/2}$ is ultimately attained for $t\gg t_{L}$ where $t_{L}\propto L^{2}/D$ is the time scale associated to the exploration of the periodic direction. Hence, we plot the VACF rescaled to its asymptotic tail [Eq. (55)] as a function of the rescaled time $\tilde{t}:=t/t_{L}$; without loss of generality we can take $t_{L}\equiv L^{2}$ and the result we obtain is the shown in Fig. 6. We observe that the data collapse to the master curve protracts from large scaled times $\tilde{t}\gg 1$ down to small times $\tilde{t}\ll 1$ upon increasing $L$. We further note that the scaling breaks down for microscopic times $t=O(\tau)$; in terms of rescaled times this implies a behavior proportional to $\tilde{t}^{-2}$ for $\tilde{t}\gg\tau/t_{L}$ and $\tilde{t}^{-3/2}$ for $\tilde{t}\gg 1$. In the limit $L\rightarrow\infty$ and $t\rightarrow\infty$ with fixed $t/t_{L}$ the data exhibit a collapse onto a single master curve whose behavior at large times is $\tilde{t}^{-3/2}$ while for short times it behaves as $\pi^{-1/2}\tilde{t}^{-2}$, these asymptotic power laws are indicated with dashed and dot-dashed black curves in Fig. 6.
Figure 6: The VACF rescaled to its asymptotic tail $nA_{L}t_{L}^{-3/2}$. The dashed black line indicates the exact long-time tail $\propto\tilde{t}^{-3/2}$, the dot-dashed corresponds to the transient tail $\propto\tilde{t}^{-2}$.
The black curve in Fig. 6 is obtained by inverting the Laplace transform $\widehat{Z}_{L}(s)$ for the unbounded model. The latter is formally obtained by taking the limit $L\rightarrow\infty$ in $\widehat{Z}_{L}(s)=[1+n(1+\mathcal{V}_{L}(0,s))]/4$, where $\mathcal{V}_{L}(0;s)=-8/\Delta_{L}(s)$ is the function already discussed in the context of the velocity relaxation, however, in this case it is evaluated for $F=0$ and $s\neq 0$. As anticipated, the limit can be performed at a formal level while its implementation is far from being trivial since propagators themselves are expressed as a sum over modes $k_{y}=2\pi q/L$ labelled by an integer running from $q=1$ to q=L1q=L-1; we refer to Appendix A for further details on lattice Green’s functions. Nonetheless, results for $L\rightarrow\infty$ are more conveniently retrieved simply by examining the solution of the unbounded model which yields [47, 54], in our notations,
@@
where $\textsf{E}[\kappa]=\int_{0}^{\pi/2}\textrm{d}\theta\left(1-\kappa\sin^{2}\theta\right)^{1/2}$ is the complete elliptic integral of the second kind [65, 66]; in this paper we adopted Mathematica’s convention for elliptic integrals. For the sake of completeness we also provide the quantity $\Delta_{L}(s)$ for the unbounded model. The matrix element of the forward-scattering matrix can expressed in terms of an elliptic integral and the result takes the form
@@
the last line holds as $s\rightarrow 0$ and is useful for the analysis of the long-time tail. The small-frequency expansion of the VACF contains smooth terms in $s$ and the first non-regular term in such an expansion is proportional to $\propto s\ln s$ with prefactor $\pi/8$. The latter yields the leading form of the long-time tail which, including the prefactor, is
@@ V.2 Diffusion coefficient and mean-square displacement away from equilibrium
where $\Xi_{L}(F)$ is known exactly for any system size $L$ and arbitrary force $F$. We emphasize that the force-dependent first-order correction $\Xi_{L}(F)$ is formally defined as
$\Xi_{L}(F)=\lim_{n\rightarrow 0}\frac{D_{x}(t\rightarrow\infty)-D_{x}^{0}(F)}{n}\,$, (61)
where $D_{x}^{0}(F)$ is the bare diffusion coefficient [Eq. (21)] and the infinite-time diffusion coefficient is found from its Laplace transform, $D_{x}(t\rightarrow\infty)=\lim_{s\rightarrow 0}s\widehat{D}(s)$, with $\widehat{D}(s)$ is obtained from the scattering formalism, as explained in Appendix B. Explicit analytic results for $\Xi_{L}(F)$ are available in the supporting Mathematica script [62]. In particular, for $F=0$ one finds $\Xi_{L}(0)=(1/4)-2/C_{L}$ and the corresponding result in equilibrium [Eq. (48)] is retrieved. In order to single out the effects due to the disorder we consider the force-induced diffusion coefficient corrected by the empty lattice ($n=0$), this subtraction entails
@@
2 0.4723 $4\sqrt{2}-4$ $1.6569$
3 0.9556 $4\sqrt{7/3}-4$ $2.1101$
4 1.2412 $\sqrt{2}+2\sqrt{3}-4 2.2$9252.2925
@@ VI Summary and conclusions
We find that – already in equilibrium – the quasi-confined lattice Lorentz model in the absence of the external force displays the intriguing feature of a dimensional crossover. We showed this property by an exact analysis of the equilibrium VACF. The latter is characterized by a transient long-time tail of the form $t^{-2}$ – reminiscent of the unbounded model –, while the long-time asymptotic behavior is a power law $\propto t^{-3/2}$. These different behaviors can be rationalized in terms of a dimensional-dependent long-time tail of the form $t^{-(d+2)/2}$. The analytic solution shows that for any finite $L$ the effective dimensionality is $d=1$ at large times while $d=2$ at intermediate times. The two regimes are separated by a size-dependent time scale $t_{L}=L^{2}/D$ which can be interpreted as the time to explore the confined direction.
@@
We then switched on the driving force. When the tracer particle is pulled a non-equilibrium stationary state is reached in which the average motion occurs with a force-dependent terminal velocity that we found analytically. The instantaneous velocity relaxes towards its terminal value via a power-law $t^{-(d+2)/2}$ decorated with an exponential time-dependence due to the force; the effective dimensionality is $d=1$ for any finite confinement $L$. As a result, the long-time tail is fragile, as it happens for the unbounded model and in such a case $d=2$. We analyzed the fluctuations of the tracer in the direction parallel to the force. For large times the variance of the displacement along the force grows linearly with time and its prefactor is a force-dependent and confinement-dependent diffusion coefficient. For large forces the confinement plays a minor role, as expected on intuitive basis since the particle hops rarely in the confined direction. On the other hand the diffusion coefficient turns out to be sensitive to the confinement in the regime of small driving. In this regime the force-dependent diffusion coefficient of the unbounded model exhibits non-analytic behavior of the form $F^{2}\ln|F|$. The occurrence of non-analyticity persists upon confinement, now with a non-analytical behavior of the form $b_{L}|F|$, with an amplitude $b_{L}$ that we obtained analytically. Although it is expected the suppression of diffusivity upon increasing the disorder, we found that this is true only when the force is smaller than a critical force $F_{c,L}$. Quite interestingly, if $F>F_{c,L}$ then the increase of disorder leads to an enhanced diffusivity. We find that $F_{c,L}$ decreases upon increasing $L$, meaning that the onset for the disorder-enhancement diffusion regime is favored by the confinement. Finally, the complete time-dependent solution shows that the behaviors at short and large times are interpolated by a transient superdiffusive behavior. We characterized this feature in terms of the local exponent of anomalous diffusion $\alpha(t)$ by showing how it increases upon increasing of the force. In this regard, especially for large forces, the confined model behaves analogously to the unbounded one and therefore the saturation to $\alpha(t)=3$ is expected as a transient phenomenon followed by diffusion at infinite times.
An important point should be made explicit. For any finite obstacle density $n>0$, there is a nonzero probability that the channel contains a blocking cluster of obstacles. As the tracer explores longer and longer distances, the probability of encountering such a cluster approaches unity; hence, at sufficiently long times the tracer becomes immobilized. Consequently, in the infinite-time limit, transport coefficients such as the diffusion coefficient vanish for any finite $n$. This implies the existence of a clogging time scale $t_{*}(n,F)$ (depending on $n$ and possibly on the drive $F$) beyond which clogging dominates. Our numerical simulations probe the pre-asymptotic regime $t\ll t_{*}$; within this time window the tracer remains mobile and no clogging is observed.